# Transfer Learning in Materials Informatics: structure-property relationships through minimal but highly informative multimodal input

### v2

@@ Abstract
In this work we propose simple, effective and computationally efficient transfer learning approaches for structure-property relation predictions in the context of materials, with highly informative input from different modalities. As materials properties stand from their electronic structure, representations are extracted directly from datasets of electronic charge density profile images using Neural Networks. We demonstrate transferability of the existing pre-trained Convolutional Neural Networks and Large Language Models knowledge to physics domain data, exploring a wide set of compositions for the regression of energetics- or structure- related properties, and the role of semantic crystallographic information in the context of multimodal approaches. We test the applicability of the CLIP multimodal model, and employ as well a training protocol for building a more interpretable and versatile stacked custom solution from different pre-trained modalities. The study offers a promising avenue for enhancing the effectiveness of descriptor identification in physical systems, shedding light on the power of multimodal transfer learning for materials property prediction, without any support of complex geometrical architectures and with potential impact on decision making for new low-cost AI methods in the field of Materials and Chemoinformatics.
@@ 3 Experiments
R2 R2 R2
#1 Val. Folds Average 0.84 0.90 0.91
Test 0.72 0.87 0.88
#2 Val. Folds Average 0.88 0.90 0.88
Test 0.89 0.82 0.80
#3 Val. Folds Average 0.77 0.90 0.92
Test 0.67 0.51 0.80
@@ Appendix A Appendix / supplemental material
Figure 3: Visualization of the samples distributions according to the three target properties, the formation energy per atom ($E_{F}$), the Fermi energy ($E_{fermi}$) and the bulk modulus ($K_{vrh}$), and to the number of constitutive atomic species ($N_{elem})$. For each violin plot, a number $n$ on the upper edge of the frame indicates the corresponding number of samples. The plot clearly indicates the majority of samples being three-species compounds. A wider (narrower) horizontal section of a single distribution instead indicates more (fewer) data points are present at that value. The white mark indicates the median, while the black bar the interquartile (IQR) range. The distributions here shown correspond already to the cleaned version of the dataset, with 592 samples.

### v3

@@ Abstract
In this work we propose simple, effective and computationally efficient transfer learning approaches for structure-property relation predictions in the context of materials, with highly informative input from different modalities. As materials properties stand from their electronic structure, representations are extracted directly from datasets of electronic charge density profile images using Neural Networks. We demonstrate transferability of the existing pre-trained Convolutional Neural Networks and Large Language Models knowledge to physics domain data, exploring a wide set of compositions for the regression of energetics- or structure- related properties, and the role of semantic crystallographic information in the context of multimodal approaches. We test the applicability of the CLIP multimodal model, and employ as well a training protocol for building a more interpretable and versatile stacked custom solution from different pre-trained modalities. The study offers a promising avenue for enhancing the effectiveness of descriptor identification in physical systems, shedding light on the power of multimodal transfer learning for materials property prediction. Instead of using the well-established GNN-based approaches, we explore the transfer learning of image- and text-based architectures, which can impact decision making for new low-cost AI methods in the field of Materials and Chemoinformatics.
@@ 3 Experiments
R2/RMSE[eV/atom] R2/RMSE[eV] R2/RMSE[GPa]
#1 Val. Folds Average 0.91/0.21 0.90/0.89 0.91/21.9
Test 0.72/0.36 0.87/0.90 0.88/24.4
#2 Val. Folds Average 0.88/0.23 0.90/0.88 0.88/24.1
Test 0.89/0.24 0.82/1.25 0.80/35.3
#3 Val. Folds Average 0.77/0.35 0.90/0.86 0.92/21.1
Test 0.67/0.35 0.51/2.07 0.80/36.5
@@ 4 Discussion, Limitations and Future Work
The Author(s) D. Massa and S. Papanikolaou were supported by the European Union Horizon 2020 research and innovation program under Grant Agreement No. 857470 and from the European Regional Development Fund under the program of the Foundation for Polish Science International Research Agenda PLUS, grant No. MAB PLUS/2018/8, and the initiative of the Ministry of Science and Higher Education ’Support for the activities of Centers of Excellence established in Poland under the Horizon 2020 program’ under agreement No. MEiN/2023/DIR/3795. Piotr Sankowski was partially supported by ERC consolidator grant TUgbOAT No. 772346 and NCN No. 2020/37/B/ST6/04179.
@@ Appendix A Appendix / supplemental material
Figure 3: Visualization of the samples distributions according to the three target properties, the formation energy per atom ($E_{F}$ in eV/atom), the Fermi energy ($E_{fermi}$ in eV) and the bulk modulus ($K_{vrh}$ in GPa), and to the number of constitutive atomic species ($N_{elem})$. For each violin plot, a number $n$ on the upper edge of the frame indicates the corresponding number of samples. The plot clearly indicates the majority of samples being three-species compounds. A wider (narrower) horizontal section of a single distribution instead indicates more (fewer) data points are present at that value. The white mark indicates the median, while the black bar the interquartile (IQR) range. The distributions here shown correspond already to the cleaned version of the dataset, with 592 samples.
@@
Figure 9: Performance of the models before and after fine-tuning.